# Algorithmic Fairness amid Social Determinants: Reflection, Characterization, and Approach

### v1

@@ A.1 Fairness Notions Based on Observational Statistics and Causal Analysis
Various notions have been proposed in the algorithmic fairness literature to characterize fairness with respect to the prediction or the prediction-based decision-making [42, 52, 29, 140], and also notions that are based on causal modeling of the data generating process [72, 67, 86, 28, 133, 36]. Recent survey papers have presented overviews on fairness notions in static settings [75, 77, 82], dynamic settings [145], and also the connection between algorithmic fairness and the literature from moral and political philosophy [125].
The type of information utilized reflects different emphases of algorithmic fairness studies. Notions based on observational statistics analyze the fairness implications in terms of the outcome of predictions or decision-making [42, 52, 29, 140, 66, 46]. Approaches that capture causal influences from the protected feature to the target variable at the individual-level [72, 67, 86, 28, 133] and the (sub-)group-level [36, 59, 83] put more emphases on the procedural aspect of algorithmic fairness inquiries, focusing on the data generating process of interest. Recent work has also proposed to address procedural fairness over all objectionable data generating components [126] according to John Rawls’s advocacy for pure procedural justice [99, 100].
@@ A.2 Detailed Comparison with Causal Fairness Approaches
Among previous algorithmic fairness approaches, causal fairness analyses are most closely related to our work since they also emphasize the role of data generating process (Section A.1). In this subsection, we provide a detailed comparison between our approach and previous works on causal fairness, in terms of the question of interest (Section A.2.1), further remarks from a purely technical perspective of causal inference (Section A.2.2), and whether or not our framework are in tension with previous causal fairness approaches (Section A.2.3).
@@ A.2.1 Question of Interest
To avoid overloading the term “counterfactual” in the causal inference literature [116, 93, 95], we use “counter-factual” (with a hyphen, as an opposite to “factual”) to denote that something does not happen in the current reality. Previous causal fairness approaches have utilized interventional [67, 86, 87, 88] and/or counterfactual [72, 28, 133] causal effects in the technical formulation, and aim to answer the following question:
@@
Based on estimating or bounding certain causal effects among variables, including the protected feature, the (predicted) outcome, and certain variables that are closely related to but not the protected feature itself, e.g., proxy variables [67], redlining attributes [144], admissible variables [108], and so on, the fairness violation is quantified in terms of causal effects between the protected feature and the (predicted) outcome. There is a reductive focus solely upon the protected feature when modeling the discrimination. For instance, it is a common practice for causal fairness notions to consider varying the value of protected feature [67, 72, 86, 87, 88, 28, 133] as the starting point. Recently, Tang et al. [126] have also proposed to consider not only edges or paths originating from the protected feature, but also all objectionable components in the data generating process, to address procedural fairness.
However, the modeling choice of “summarizing” discrimination only through edges/paths originating from protected feature, or solely among individual-level variables, falls short of the need to capture procedural unfairness and structural injustice. The characteristics of the environment and the context that individuals operate in typically do not correspond to individual-level attributes, and are not considered in previous literature. Different from causal fairness approaches, our approach calls for explicit incorporations of the influence of contextual environments, and aims to address the following question:
@@
A.2.2 Further Remarks from the Technical Perspective of Causal Inference
From a purely technical point of view, it is not trivial to incorporate social determinants as just another set of variables into existing causal fairness approaches.
To begin with, social determinants necessitates community-level considerations that go beyond individual-level comparisons, which are typically sufficient when considering sensitive attributes alone. For instance, for (path-specific) causal fairness notions, which characterize causal effects at the individual level, frame the causal effect originating solely from sensitive attributes. The contrast is between worlds in which sensitive attributes of an individual were different. However, social determinants, which may be community-level attributes, are essentially formed by the people and environment around the specific individual of interest. This indicates the existence an additional structure among individuals and/or environment in that community, which are formed exactly by these individuals collectively. In other words, the individual-level DAG specifications do not readily capture such community-level variables, whose values are closely related to the composition of individuals in that community and/or environment.
Furthermore, whenever controlling for, or performing intervention on, social determinants, both this individual and the contexts around them are subject to downstream effects. For instance, because of the total amount of resources available in the community formed by this set of individuals, “intervening on” a community-level variable (e.g., a summary statistics) of an individual necessitates accounting for the redistribution of resource that involves all individuals in the community. This is very different from current causal fairness analyses (e.g., the controlling for confounding when establishing causal effect identification in causal inference), where after controlling for individual-level variables, the effect of which does not extend beyond the individual of interest. Previous works also consider aggregation of individual-level causal effects among subgroups [36, 59, 83], but the starting point remains individual-level causal modeling [72, 67, 86, 28, 133].
Therefore, from a purely technical point of view, incorporating social determinants involves nontrivial developments of new fairness notions (that dynamically capture changes in contexts and environments), data collection and processing schemes, and mitigation strategies.
A.2.3 No Conflict in Principle with Causal Fairness
@@
However, our proposal extends the scope of consideration beyond sensitive variables, and explicitly incorporates the influence of contextual environments. For instance, when operationalizing our proposal, we do not drop relevant variables, e.g., the Address of an individual, which is often omitted in previous literature [67, 72, 86, 28, 133, 80, 40]. Furthermore, the findings of our analyses suggest that we should utilize all information available, and furthermore, actively look for and develop better measurements for social determinants, so that we can better understand and address structural injustice. Future works naturally include the development of causal effect estimands that incorporate both sensitive attributes and social determinants, and our proposal and previous causal fairness approaches can be used in conjunction to achieve the goal.
@@ A.3.1 The Uses of “Structure” in Related Disciplines
The term “structure” and “structural” are utilized in different ways by related disciplines. For the literature of causal learning and reasoning, the term “structure” and “structural” are often used to describe how causal structures look like among variables of interest [116, 93, 95, 54], e.g., in terms of causal graphs and/or structural equation models (SEMs). For the literature of structural justice and social determinants, the term “structural” is used to denote the systemic ways in which society is organized, e.g., through policies, laws, and social norms, that perpetuate discrimination and animus towards certain groups [20, 114, 128, 134, 104, 6, 135]. There are interests in the social determinants of health literature to use DAGs as a tool for illustrative purposes, abstracting key concepts or areas that are interrelated at a high level, and modeling the mechanism through which structural forms of discriminations get realized (racism, sexism, etc.) [104, 135].
@@ A.3.2 Conceptual Distinctions Between “Unfairness” and “Discrimination”
The term “discrimination” refers to actions, practices, or policies that are based on the (perceived) social group membership of those affected. Standard accounts mandate that these groups are socially salient, i.e., they must significantly shape interactions within important social contexts [73, 96, 4] while recent works have challenged the social salience requirement [43]. The term unfairness is typically understood as the broader concept, which encompassing any violation of principles of justice or proper treatment [5, 85]. In algorithmic fairness literature, existing fairness inquiries (including observational and causal ones) tend to gravitate towards quantifying discrimination. Meanwhile, achieving fairness (through addressing social determinants) receives less attention compared to enforcing fairness (through addressing sensitive attributes).
@@ A.4 Common Presence of Social Determinants
In terms of the influence of environments on individual’s health, previous literature has considered how environmental hazards disproportionately affect low-income populations and communities of color [131], how indoor air pollution affects women and children in low-income regions [78], and the structural implications of social determinants on how society should be organized [104, 135]. More broadly, a review on economic research has also been conducted to show how environmental changes impact public health in both developed and developing countries [102].
@@
In terms of the influence of environments on individual’s educational attainments, previous literature has considered how the quality of schools and the availability of educational resources affect students’ academic performance [31, 32], how the family and neighborhood environments influence education [60], and implications of various affirmative-action policies (usually under different names) across countries with different histories and cultures [115].
@@
In terms of the influence of environments on individual’s employment opportunities, previous literature has considered the relation between the employment of residents and the rationalization and optimization level of region’s industrial structures [19, 98], the psychological perspective of (e.g., influence from collective values of community) job search behaviors [129], and how the employment rate of residents is influenced by job quality [56].
@@ Appendix B Background Information of Admission Strategies & Proofs of Theoretical Results
In this section, we provide background information of admission strategies and present proofs of our theoretical results.
@@ B.1 Background Information of Admission Strategies
The quota-based admission is a type of affirmative-action admission strategy that sets specific limits on the number of admissions for applicants from different demographic backgrounds. This admission strategy was originally designed to rectify historical injustice by directly setting aside admission quotas to increase the representation of URM students. However, due to the rigid nature of the quota-based mechanism, this admission strategy has been controversial and addressed by the U.S. Supreme Court in the landmark case University of California Regents v. Bakke (1978) [118]. It was held that the use of strict racial quotas in college admission was unconstitutional, and was reaffirmed in another landmark case Grutter v. Bollinger (2003) [120].
@@
Holistic review with plus factors is another type of affirmative-action admission strategy, involving consideration of multiple factors that together define each individual applicant. The key element of this process is the use of plus factors, where certain characteristics, for instance, race and ethnic group, are given additional weight to promote diversity in the student body and rectify historical disadvantages. This approach was upheld by the U.S. Supreme Court in Grutter v. Bollinger (2003) [120], but was overruled in recent decisions for Students for Fair Admissions (SFFA) v. Harvard & UNC (2023) [121, 122], effectively banning race-conscious admissions.
For holistic review with plus factors, we model its affirmative-action emphasis on the URM group through a distribution shift, i.e., from the original scale to the plus-factor scale, instead of an automatic awarding of points for each URM applicant. Our modeling choice is for the purpose of avoiding the introduction of rigid and mechanical characteristics to the process, as was addressed in Gratz v. Bollinger (2003) [119].
@@
The top-percentage plans are college admission policies that guarantee admission to students who graduate in a certain top percentage of their high school classes. The top-percentage plans are generally not considered traditional affirmative-action admission strategies. Instead, these policies are race-neutral alternatives aiming to promote diversity by drawing students from a wide range of schools with different socioeconomic and geographic backgrounds, without explicitly considering race. A prominent example is the University of Texas’s Top 10% Rule, which guarantees admission to students in the top 10% of their class. Another is the Eligibility in the Local Context (ELC) program of University of California, which was introduced after the 1996 California Proposition 209 banned the use of race, ethnicity, and gender in public university admissions in California.
@@
Theorem (Quota-Based Admission Incurs Unfairness w.r.t. Non-URM in Poor Region).
@@
$\max_{q\in[0,\infty)}\frac{F^{(\mathrm{rich})}(q)}{F^{(\mathrm{poor})}(q)}\geq\frac{(n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})})\eta_{\mathrm{quota}}}{(n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})})(1-\eta_{\mathrm{quota}})+(n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})})}\leavevmode\nobreak\ \raisebox{-7.74997pt}{.}$ (B.1)
@@
  • LHS equals to RHS of Equation (B.4): since $F^{(\mathrm{rich})}$ dominates $F^{(\mathrm{poor})}$ (Assumption 3.3), we have $q_{a}^{\text{(}poor)}>q_{a}^{\text{(}rich)}$, i.e., among URM applicants, the threshold for the raw score in the poor region is lower than that for the rich region.
@@
where (i) results from Equations B.4 and B.5, and (ii) follows Equation (B.4).
@@
$\max_{q\in[0,\infty)}\frac{F^{(\mathrm{rich})}(q)}{F^{(\mathrm{poor})}(q)}\geq\frac{(n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})})\eta_{\mathrm{quota}}}{(n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})})(1-\eta_{\mathrm{quota}})+(n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})})}\leavevmode\nobreak\ \raisebox{-7.74997pt}{.}$
B.3 Proof of Theorem 3.6 in Section 3.3
Theorem (Holistic Review with Plus Factors Benefits URM in Rich Region More).
Under Assumptions 3.13.4, let us denote with $\eta_{\dagger}<1$ the multiplicative coefficient on the scale parameter of Gamma distributions for URM applicants’ academic index scores, such that the perceived scores of URM applicants shift more probability density towards the high-score end. Let us denote with $q_{o}$ the default threshold for selective admission, and with $q_{\dagger}$ the threshold if the admission procedure is a holistic review with plus factors. Further assume that region-specific shape parameters satisfy $k^{(\mathrm{poor})}=k^{(\mathrm{rich})}=k_{o}$. Then, the increase in the probability of selective admission for URM applicants from the rich region, is larger than that for URM applicants from the poor region:

### v2

@@ A.1 Fairness Notions Based on Observational Statistics and Causal Analysis
Various notions have been proposed in the algorithmic fairness literature to characterize fairness with respect to the prediction or the prediction-based decision-making (Dwork et al., 2012; Hardt et al., 2016; Chouldechova, 2017; Zafar et al., 2017), and also notions that are based on causal modeling of the data generating process (Kusner et al., 2017; Kilbertus et al., 2017; Nabi & Shpitser, 2018; Chiappa, 2019; Wu et al., 2019; Coston et al., 2020). Recent survey papers have presented overviews on fairness notions in static settings (Loftus et al., 2018; Makhlouf et al., 2020; Mehrabi et al., 2021), dynamic settings (Zhang & Liu, 2021), and also the connection between algorithmic fairness and the literature from moral and political philosophy (Tang et al., 2023).
The type of information utilized reflects different emphases of algorithmic fairness studies. Notions based on observational statistics analyze the fairness implications in terms of the outcome of predictions or decision-making (Dwork et al., 2012; Hardt et al., 2016; Chouldechova, 2017; Zafar et al., 2017; Kearns et al., 2018; Foulds et al., 2020). Approaches that capture causal influences from the protected feature to the target variable at the individual-level (Kusner et al., 2017; Kilbertus et al., 2017; Nabi & Shpitser, 2018; Chiappa, 2019; Wu et al., 2019) and the (sub-)group-level (Coston et al., 2020; Imai & Jiang, 2020; Mishler et al., 2021) put more emphases on the procedural aspect of algorithmic fairness inquiries, focusing on the data generating process of interest. Recent work has also proposed to address procedural fairness over all objectionable data generating components (Tang et al., 2024) according to John Rawls’s advocacy for pure procedural justice (Rawls, 1971, 2001).
@@ A.2 Detailed Comparison with Causal Fairness Approaches
In this subsection, we provide a detailed comparison between our position and previous works on causal fairness, in terms of the question of interest (Section A.2.1), and whether or not our framework are in tension with previous causal fairness approaches (Section A.2.2).
@@ A.2.1 Question of Interest
To avoid overloading the term “counterfactual” in the causal inference literature (Spirtes et al., 1993; Pearl, 2000; Peters et al., 2017), we use “counter-factual” (with a hyphen, as an opposite to “factual”) to denote that something does not happen in the current reality. Previous causal fairness approaches have utilized interventional (Kilbertus et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019, 2022) and/or counterfactual (Kusner et al., 2017; Chiappa, 2019; Wu et al., 2019) causal effects in the technical formulation, and aim to answer the following question:
@@
Based on estimating or bounding certain causal effects among variables, including the protected feature, the (predicted) outcome, and certain variables that are closely related to but not the protected feature itself, e.g., proxy variables (Kilbertus et al., 2017), redlining attributes (Zhang et al., 2017), admissible variables (Salimi et al., 2019), and so on, the fairness violation is quantified in terms of causal effects between the protected feature and the (predicted) outcome. There is a reductive focus solely upon the protected feature when modeling the discrimination. For instance, it is a common practice for causal fairness notions to consider varying the value of protected feature (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019, 2022; Chiappa, 2019; Wu et al., 2019) as the starting point. Recently, Tang et al. (2024) have also proposed to consider not only edges or paths originating from the protected feature, but also all objectionable components in the data generating process, to address procedural fairness.
However, the modeling choice of “summarizing” discrimination only through edges/paths originating from protected feature, or solely among individual-level variables, falls short of the need to capture structural injustice. The characteristics of the environment and the context that individuals operate in typically do not correspond to individual-level attributes, and are not considered in previous literature. Different from causal fairness approaches, our position calls for explicit incorporations of the influence of contextual environments, and aims to address questions like:
@@
A.2.2 No Conflict in Principle with Causal Fairness
@@
However, our proposal extends the scope of consideration beyond sensitive variables, and explicitly incorporates the influence of contextual environments. For instance, when operationalizing our proposal, we do not drop relevant variables, e.g., the Address of an individual, which is often omitted in previous literature (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Chiappa, 2019; Wu et al., 2019; Mary et al., 2019; Ding et al., 2021). Furthermore, the findings of our analyses suggest that we should utilize all information available, and furthermore, actively look for and develop better measurements for social determinants, so that we can better understand and address structural injustice. Future works naturally include the development of causal effect estimands that incorporate both sensitive attributes and social determinants, and our proposal and previous causal fairness approaches can be used in conjunction to achieve the goal.
@@ A.3.1 The Uses of “Structure” in Related Disciplines
The term “structure” and “structural” are utilized in different ways by related disciplines. For the literature of causal learning and reasoning, the term “structure” and “structural” are often used to describe how causal structures look like among variables of interest (Spirtes et al., 1993; Pearl, 2000; Peters et al., 2017; Hernán & Robins, 2020), e.g., in terms of causal graphs and/or structural equation models (SEMs). For the literature of structural justice and social determinants, the term “structural” is used to denote the systemic ways in which society is organized, e.g., through policies, laws, and social norms, that perpetuate discrimination and animus towards certain groups (Carmichael et al., 1967; Sowell, 1972; Tilly, 1998; Yearby, 2018; Robinson et al., 2020; Alexander, 2020; Yearby et al., 2022). There are interests in the social determinants of health literature to use DAGs as a tool for illustrative purposes, abstracting key concepts or areas that are interrelated at a high level, and modeling the mechanism through which structural forms of discriminations get realized (racism, sexism, etc.) (Robinson et al., 2020; Yearby et al., 2022).
@@ A.3.2 Conceptual Distinctions Between “Unfairness” and “Discrimination”
The term “discrimination” refers to actions, practices, or policies that are based on the (perceived) social group membership of those affected. Standard accounts mandate that these groups are socially salient, i.e., they must significantly shape interactions within important social contexts (Lippert-Rasmussen, 2006; Porat, 2005; Al Ramiah et al., 2010) while recent works have challenged the social salience requirement (Eidelson, 2015). The term unfairness is typically understood as the broader concept, which encompassing any violation of principles of justice or proper treatment (Alexander, 1992; Moreau, 2020). In algorithmic fairness literature, existing fairness inquiries (including observational and causal ones) tend to gravitate towards quantifying discrimination. Meanwhile, achieving fairness as structural justice (through addressing social determinants) receives less attention compared to enforcing fairness as non-discrimination (through addressing sensitive attributes).
@@ A.4 Common Presence of Social Determinants
In terms of the influence of environments on individual’s health, previous literature has considered how environmental hazards disproportionately affect low-income populations and communities of color (Warren et al., 2002), how indoor air pollution affects women and children in low-income regions (Manisalidis et al., 2020), and the structural implications of social determinants on how society should be organized (Robinson et al., 2020; Yearby et al., 2022). The time consistency and (sub-)population performance disparity of a mortality predictor has been evaluated in the context of advance care planning (Handler et al., 2023). More broadly, a review on economic research has also been conducted to show how environmental changes impact public health in both developed and developing countries (Remoundou & Koundouri, 2009).
@@
In terms of the influence of environments on individual’s educational attainments, previous literature has considered how the quality of schools and the availability of educational resources affect students’ academic performance (Coleman, 1968, 1988), how the family and neighborhood environments influence education (Jencks, 1972), and implications of various affirmative-action policies (usually under different names) across countries with different histories and cultures (Sowell, 2004).
@@
In terms of the influence of environments on individual’s employment opportunities, previous literature has considered the relation between the employment of residents and the rationalization and optimization level of region’s industrial structures (Cao et al., 2017; Qin et al., 2022), the psychological perspective of (e.g., influence from collective values of community) job search behaviors (van Hooft et al., 2021), and how the employment rate of residents is influenced by job quality (Howell & Kalleberg, 2019).
@@ Appendix B Background Information of Admission Strategies & Proofs of Theoretical Results
In this section, we provide background information of admission strategies, present additional theoretical results and the proofs.
@@ B.1 Background Information of Admission Strategies
The quota-based admission is a type of affirmative-action admission strategy that sets specific limits on the number of admissions for applicants from different demographic backgrounds. This admission strategy was originally designed to rectify historical injustice by directly setting aside admission quotas to increase the representation of URM students. However, due to the rigid nature of the quota-based mechanism, this admission strategy has been controversial and addressed by the U.S. Supreme Court in the landmark case University of California Regents v. Bakke (1978) (Supreme Court, 1978). It was held that the use of strict racial quotas in college admission was unconstitutional, and was reaffirmed in another landmark case Grutter v. Bollinger (2003) (Supreme Court, 2003b).
@@
Holistic review with plus factors is another type of affirmative-action admission strategy, involving consideration of multiple factors that together define each individual applicant. The key element of this process is the use of plus factors, where certain characteristics, for instance, race and ethnic group, are given additional weight to promote diversity in the student body and rectify historical disadvantages. This approach was upheld by the U.S. Supreme Court in Grutter v. Bollinger (2003) (Supreme Court, 2003b), but was overruled in recent decisions for Students for Fair Admissions (SFFA) v. Harvard & UNC (2023) (Supreme Court, 2023a, b), effectively banning race-conscious admissions.
For holistic review with plus factors, we model its affirmative-action emphasis on the URM group through a distribution shift, i.e., from the original scale to the plus-factor scale, instead of an automatic awarding of points for each URM applicant. Our modeling choice is for the purpose of avoiding the introduction of rigid and mechanical characteristics to the process, as was addressed in Gratz v. Bollinger (2003) (Supreme Court, 2003a).
@@
The top-percentage plans are college admission policies that guarantee admission to students who graduate in a certain top percentage of their high school classes. The top-percentage plans are generally not considered traditional affirmative-action admission strategies. Instead, these policies are race-neutral alternatives aiming to promote diversity by drawing students from a wide range of schools with different socio-economic and geographic backgrounds, without explicitly considering race. A prominent example is the University of Texas’s Top 10% Rule, which guarantees admission to students in the top 10% of their class. Another is the Eligibility in the Local Context (ELC) program of University of California, which was introduced after the 1996 California Proposition 209 banned the use of race, ethnicity, and gender in public university admissions in California.
@@
Theorem.
@@
$\max_{q\in[0,\infty)}\frac{F^{(\mathrm{rich})}(q)}{F^{(\mathrm{poor})}(q)}\geq\frac{(n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})})\eta_{\mathrm{quota}}}{(n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})})(1-\eta_{\mathrm{quota}})+(n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})})}\penalty 10000\ \raisebox{-8.61108pt}{.}$ (B.1)
@@
  • LHS equals to RHS of Equation (B.4): since $F^{(\mathrm{rich})}$ dominates $F^{(\mathrm{poor})}$ (Assumption B.1), we have $q_{a}^{\text{(}poor)}>q_{a}^{\text{(}rich)}$, i.e., among URM applicants, the threshold for the raw score in the poor region is lower than that for the rich region.
@@
where (i) results from Equation (B.4) and Equation (B.5), and (ii) follows Equation (B.4).
@@
$\max_{q\in[0,\infty)}\frac{F^{(\mathrm{rich})}(q)}{F^{(\mathrm{poor})}(q)}\geq\frac{(n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})})\eta_{\mathrm{quota}}}{(n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})})(1-\eta_{\mathrm{quota}})+(n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})})}\penalty 10000\ \raisebox{-8.61108pt}{.}$
(a) Quota-based admission
(b) Holistic review with plus factors
(c) Top-percentage plan
B.3 Additional Theoretical Result Related to Holistic Review with Plus Factors
For the purpose of deriving a closed-from characterization, we introduce an additional quantitative assumption:
Assumption B.1 (Gamma Parameterization of Academic Preparedness Distribution).
[…]
In Assumption B.1, the conversion of the score maps the domain of values $S_{\text{MIN}},S_{\text{MAX}}]$ (higher score $S$ is more competitive) to $0,\infty)$, where the closer to 0 the converted score $Q$, the more competitive. We selected Gamma distributions for their flexibility in capturing the one-sided skew commonly observed in academic performance distributions, consistent with prior educational assessment research (Campos, 2013; Sari et al., 2025).
Putting aside the evolving jurisprudence (Supreme Court, 2003b, 2023a, 2023b), we aim to precisely characterize holistic review in terms of its implications on the distribution of benefits and burdens among individuals, when allocating the limited spots in selective college admissions. When taking into account of social determinants signified by Region, we show that holistic review with plus factors may benefit applicants from better-off areas more than those from less well-off areas:
Theorem B.2 (Holistic Review with Plus Factors Benefits URM in Rich Region More).
Under Assumptions 4.14.4 and also B.1, let us denote with $\eta_{\dagger}<1$ the multiplicative coefficient on the scale parameter of Gamma distributions for URM applicants’ academic index scores, such that the perceived scores of URM applicants shift more probability density towards the high-score end. Let us denote with $q_{o}$ the default threshold for selective admission, and with $q_{\dagger}$ the threshold if the admission procedure is a holistic review with plus factors. Further assume that region-specific shape parameters satisfy $k^{(\mathrm{poor})}=k^{(\mathrm{rich})}=k_{o}$. Then, the increase in the probability of selective admission for URM applicants from the rich region, is larger than that for URM applicants from the poor region: